### v1

# Bridging the Gap Between LLMs and Human Intentions: Progresses and Challenges in Instruction Understanding, Intention Reasoning, and Reliable Generation

Abstract
Large language models (LLMs) have demonstrated exceptional capabilities in understanding and generation. However, when interacting with human instructions in real-world scenarios, LLMs still face significant challenges, particularly in accurately capturing and comprehending human instructions and intentions. This paper focuses on three challenges in LLM-based text generation tasks: instruction understanding, intention reasoning, and reliable generation. Regarding human complex instruction, LLMs have deficiencies in understanding long contexts and instructions in multi-round conversations. For intention reasoning, LLMs may have inconsistent command reasoning, difficulty reasoning about commands containing incorrect information, difficulty understanding user ambiguous language commands, and a weak understanding of user intention in commands. Besides, In terms of reliable generation, LLMs may have unstable generated content and unethical generation. To this end, we classify and analyze the performance of LLMs in challenging scenarios and conduct a comprehensive evaluation of existing solutions. Furthermore, we introduce benchmarks and categorize them based on the aforementioned three core challenges. Finally, we explore potential directions for future research to enhance the reliability and adaptability of LLMs in real-world applications.
1 Introduction
Rapid advancements with the development of large language models (LLMs) have been experienced in the field of artificial intelligence. These models, built upon massive amounts of data and extensive computing resources, have shown impressive capabilities in understanding and generating human language. Recent advancements in LLMs, including the use of scaling laws Kaplan et al. (2020), supervised fine-tuning (SFT) Wu et al. (2021), and reinforcement learning with human feedback (RLHF) Ouyang et al. (2022), have propelled these models to new heights. Researchers have explored innovative strategies like chain-of-thought reasoning (COT) Wei et al. (2022), aiming to enhance their performance in processing and generating accurate responses. However, they still struggle with more complex interactions, especially when the input data is ambiguous, incomplete, or inconsistent. Despite improvements, issues such as content hallucination Li et al. (2023) and logical misinterpretations remain prevalent. Consequently, while LLMs show promise, they are far from flawless and require further refinement to address the challenges posed by more unpredictable and complex human instructions as follows.
Figure 1: Example of large language model generation.
I. Challenge of Instruction Understanding. One of the most pressing challenges that LLMs face is instruction understanding as Figure 1(a), particularly when the user input involves complex or multi-step instructions. While models have improved in parsing relatively simple queries, they continue to encounter significant difficulties when dealing with long, context-rich instructions or when instructions are spread across multiple conversational turns. LLMs often fail to grasp subtle nuances or interpret implicit meanings embedded within the text, which leads to inaccurate or incomplete responses. Existing approaches to instruction understanding have introduced techniques like optimizing the model’s parsing abilities Teng et al. (2024), and context-aware optimization Sun et al. (2024). While these methods show promise, they often fall short when addressing the complexities and ambiguities present in real-world instructions.
II. Challenge of Intention Reasoning. Another critical area is intention reasoning as illustrated in Figure 1(b), where they struggle to align the generated responses with the user’s underlying intention. Ambiguities in language, conflicting instructions, and implicit requirements often result in models producing outputs that diverge from the user’s expectations. LLMs also face difficulties when instructions are inconsistent or contain incorrect information, which challenges the model’s ability to make accurate inferences. Various strategies, including retrieval-enhanced generation and fine-tuning techniques, have been proposed to enhance reasoning capabilities, enabling the models to better handle inconsistent or incomplete instructions. However, these methods often introduce new challenges related to bias and the inability to fully resolve conflicts in user input, further complicating the alignment between generated content and user expectations.
III. Challenge of Reliable Generation. The final major challenge is the reliable generation, which pertains to the accuracy, ethical considerations, and stability of the content they produce, such as Figure 1(c). While LLMs are generally capable of generating coherent and contextually relevant outputs, they sometimes exhibit instability, generating content that is factually incorrect, logically inconsistent, or ethically questionable. This challenge is exacerbated by the model’s inability to recognize uncertainty, which can lead to overconfident but inaccurate outputs. Recent efforts to address this issue involve techniques like uncertainty-aware fine-tuning and using external tools to evaluate output credibility. However, these approaches struggle to provide a comprehensive and reliable solution, especially in complex or dynamic contexts.
{forest}
forked edges, for tree= grow=east, reversed=true, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, anchor=center , [ Gap Between LLMs and Human Intention, ver, color=hidden-draw, fill=mygray!30, text=black [ Instruction
Understanding (§2), color=myblue!100, fill=myblue!80, text width=6em, text=black [ Long-text
Comprehension (§2.1), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ Information
Focusing, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) Attention Sparsification Beltagy et al. (2020)
2) Attention OptimizationChen et al. (2024)
3) Position independent Training He et al. (2024a) , understanding_leaf, text width=26.5em ] ] [ Multipath
Optimization, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) SFT and RL Zhang et al. (2024b)
2) Retrieval-Augmented Li et al. (2024b)
3) Recurrent Sequence Optimization Gu and Dao (2023) , understanding_leaf, text width=26.5em ] ] ] [ Multi-turn Conversation
Handling (§2.2), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ 1) SFT Teng et al. (2024); Sun et al. (2024), 2) RL Zhou et al. (2024) , understanding_leaf, text width=32em ] ] ] [ Intention
Reasoning (§3), color=mygreen!80, fill=mygreen!80, text width=6em, text=black [ Inconsistent Instruction
Resolution (§3.1), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Knowledge
Updating, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ SituatedQA Zhang and Choi (2021), CDConv Zheng et al. (2022) , reasoing_leaf, text width=26.5em ] ] [ Confidence
Calibration, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ CD2 Jin et al. (2024), MacNoise Hong et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] [ Misinformation
Reasoning (§3.2), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ CKL Jang et al. (2022), StreamingQA Liska et al. (2022), BIPIA Yi et al. (2023),
CAR Weller et al. (2024) , reasoing_leaf, text width=32em ] ] [ Fuzzy Language
Interpretation (§3.3), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Folkscope Yu et al. (2023), Miko Lu et al. (2024), ATC Deng et al. (2023) , reasoing_leaf, text width=32em ] ] [ Intention
Clarification
Failure (§3.4) , color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ DeepSeek-R1 Guo et al. (2025), S1 Muennighoff et al. (2025),
SoulChat Chen et al. (2023), MoChat Mo et al. (2024) , reasoing_leaf, text width=32em ] ] ] [ Reliable
Generation (§4), color=lightkhaki!80, fill=lightkhaki!80, text width=6em, text=black [ Response Stability (§4.1), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Fine-tuning
LLMs, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ EDL Sensoy et al. (2018), DER Amini et al. (2020),
ConformalFactuality Mohri and Hashimoto (2024), LUQ Zhang et al. (2024a) , reliable_leaf, text width=26.5em ] ] [ External Tools, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ SUE Liu et al. (2024), CalibrateMath Lin et al. (2022) , reliable_leaf, text width=26.5em ] ] ] [ Alignment (§4.2), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Data Cleaning
and Curation, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ Stochastic Parrots Bender et al. (2021) , DeepSoftDebias Rakshit et al. , reliable_leaf, text width=26.5em ] ] [ RL-based
Alignment , color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ PPO Ouyang et al. (2022), DPO Zeng et al. (2024), GRPO Shao et al. (2024),
RLAIF Lee et al. (2024) , reliable_leaf, text width=26.5em ] ] [ In-context
Alignment, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ URIAL Lin et al. (2024), ICA Huang et al. (2024a) , reliable_leaf, text width=26.5em ] ] ] ] ]
Figure 2: Challenges and existing solutions between LLMs and Human Intentions.
Present Survey.
Facing these challenges, there is an increasing need for focused research on LLMs and their interaction with human instructions and intentions. This paper systematically analyzes LLMs’ performance in processing human instructions, highlighting three key areas: user instruction understanding, intention comprehension and reasoning, and reliable content generation. While existing review papers address model training, fine-tuning, and specific aspects of LLMs’ capabilities Lou et al. (2024); Plaat et al. (2024); Huang et al. (2024b), our focus is on the LLMs’ ability to understand and reason about user intentions. Specifically, we explore how well LLMs understand user input, reason about the user’s intention, and generate content that aligns with human intentions, minimizing hallucinations and ensuring credibility and ethical consistency.
Comparison with Previous Surveys.
While the gap between human intention and LLMs is a core challenge in generative AI, many studies focus on specific aspects of the issue, lacking a comprehensive overview. These works offer valuable insights but do not provide a systematic summary of the field. Lou et al. Lou et al. (2024) primarily address instruction following challenges in LLMs without delving into the reasoning capabilities for complex user instructions. Gao et al. analyze the four stages of human-machine LLM interaction (planning, facilitation, iteration, and testing) but overlook LLM’s understanding of user instructions. Xu et al. Xu et al. (2024b) examine the impact of various memory conflicts on LLM-generated content credibility and performance, yet do not consider reasoning or intention comprehension. Plaat et al. Plaat et al. (2024) focus on LLM Reasoning for basic mathematical problems, without exploring its applicability to broader fields. Shorinwa et al. Shorinwa et al. (2024) provide an initial analysis of LLMs uncertainty quantification, but exclude user input instructions. In contrast, our survey offers a more comprehensive perspective, with a unique classification and systematic analysis of instruction processing, addressing current solutions to key challenges.
Survey Organization.
As in Figure 2, we begin by exploring the capability of user instruction understanding (§2). Next, we analyze intention comprehension and reasoning, focusing on how models infer implicit intentions, incorporate contextual information for logical reasoning, and address inconsistencies or incomplete instructions (§3). We then examine reliable content generation, assessing the quality and credibility of model-generated outputs (§4). Following this, we review benchmarks that address these core challenges (§5). Finally, we propose potential research directions (§6) and summarize the key findings (§7).
Figure 3: Common understanding issues in LLMs include: (A) Remote Information Failure (§2.1), where the model forgets relevant information over long distances in long context (A.1). (B) Incorrect Relevance Judgment(§2.2), such as the model incorrectly associates wrong content (B.1-2.nd) from the previous turn (B.1-1.st).
2 Instruction Understanding
LLMs excel at single-turn dialogues, but struggle to understand multi-turn dialogues and long-contexts, which are commonly used by users. LLMs may forget prior information, be influenced by irrelevant data, and overlook key inputs.
2.1 Long-Text Comprehension
Challenges remain in understanding long context and this paper classifies the factors into three categories as follows:1) Information Sparsity and Redundancy. Long texts often contain redundant or irrelevant information that can obscure the task-relevant content, leading to difficulties in information extraction and potential hallucinations. 2) Remote Information Failure (Figure 3 A.1). Long contexts may cause models to forget relevant information that is distant within the text. Additionally, links between remote information across paragraphs or sentences can be difficult for models to identify, diminishing their understanding of contextual connections. 3) Attention Dilution. As context length increases, the model’s attention mechanism faces greater computational demands and struggles to assign appropriate weights to each token, making it harder to prioritize key information, particularly with complex, multi-level relationships in longer texts. This paper classifies the existing solutions into the following two categories:
Information Focusing.
Improving LLM’s ability to focus on important information in long texts involves several methods: 1) Sparsifying attention to concentrate on critical information Beltagy et al. (2020). 2) Optimizing attention to minimize redundancy and emphasize core content Chen et al. (2024). 3) Training with location-independent tasks to enhance the ability to search and react to relevant information in long contexts He et al. (2024a).
Multipath Optimization.
Various methods can enhance LLMs on long-context tasks: 1) Pre-training with extended context windows and reinforcement learning for fine-tuning to optimize long-context understanding Zhang et al. (2024b). 2) Combining retrieval-based models with generative models on long-context tasks Li et al. (2024b). 3) Leveraging cyclic sequence models’ linear scaling property for better inference efficiency Gu and Dao (2023).
2.2 Multi-Turn Conversation Handling
This paper categorizes the challenges faced by existing LLMs when understanding multi-turn conversations into three categories, as follows: 1) Capability Weakening. Current supervised instruction fine-tuning (SIFT) and RLHF may even impair multi-turn capabilitiesWang et al. (2024), with models struggling on complex reasoning tasks that span multiple rounds, such as those requiring evidence collection and conclusions Banatt et al. (2024). Additionally, multi-turn dialogs increase the vulnerability of LLMs to adversarial attacks, where malicious users can mask harmful intentions across multiple rounds, leading to the generation of misleading or harmful content Agarwal et al. (2024). 2) Error Propagation. Instruction comprehension errors accumulate across rounds, leading to an escalating failure rate in subsequent responses He et al. (2024b), which may snowball into larger issues such as biased or incorrect outputs Fan et al. (2024). 3) Incorrect Relevance Judgment (Figure 3 B.1) LLMs often struggle to identify relevant content in multi-turn dialogs, failing to properly link content from previous rounds or to discern ellipsis and implicit meaning inherent in user commands Sun et al. (2024).
To solve above challenges, This paper categorizes existing solutions into two types: supervised fine-tuning methods using multi-turn dialogue data, enhanced by techniques like optimized instruction parsing Teng et al. (2024) and context-aware preference strategies Sun et al. (2024); and reinforcement learning methods tailored for multi-turn dialogue, with improvements such as hierarchical reinforcement learning Zhou et al. (2024).
3 Intention Reasoning
User instructions often lack clarity due to language ambiguities. While humans can infer intention, LLMs struggle with misinterpreting ambiguous inputs, leading to errors. We explore causes and solutions for intention errors, focusing on inconsistent instructions, misinformation, fuzzy language, and intention clarification.
Figure 4: Common reasoning issues in LLMs include: (A) Inconsistent Instruction Reasoning (§3.1), where LLMs fail to detect conflicting inputs (A.1) or overlooking logical inconsistencies (A.2), (B) Misinformation Reasoning (§3.2), caused by temporal misalignment leading to outdated responses (B.1) or data contamination resulting in misleading outputs (B.2). (C) Fuzzy Language Interpretation (§3.3), where the model relies on biases for fuzzy queries (C.1) or defaults to a response without seeking clarification (C.2). and (D) Intention Clarification Failure (§3.4) , where it misinterprets sarcasm (D.1) or ignores prior emotional context (D.2).
3.1 Inconsistent Instruction Reasoning
In natural language communication, humans easily identify inconsistencies using context and prior knowledge, whereas LLMs struggle, often accept contradictory inputs, and generate unreliable answers. This phenomenon has been observed across multiple question-answering generation tasks Li et al. (2024a); Zheng et al. (2022), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Ignoring input errors (Figure 4 A.1). The model ignores the input errors and gives an answer, resulting in the model assigning the same weight to each context given by the user, which in turn affects the generation of the answer. 2) Inability to detect user inconsistencies (Figure 4 A.2). In the premise that the model has learned the knowledge, the model still has difficulty detecting user inconsistencies. To address inconsistent instruction reasoning issues, existing solutions primarily adopt the following two approaches:
Knowledge Updating.
SituatedQA Zhang and Choi (2021) attempts to enhance model performance by updating the knowledge base. Additionally, CDConv Zheng et al. (2022) simulates common user behaviors to trigger chatbots through an automated dialogue generation method, generating contradictions for training purposes.
Confidence Calibration.
Given the high cost of data annotation and model fine-tuning, some researchers have sought alternative approaches by introducing additional processing techniques. CD2 Jin et al. (2024) maximizes probabilistic output and calibrates model confidence under knowledge conflicts using conflict decoupling and comparison decoding methods. MacNoise Hong et al. (2024) enhances contradiction retrieval in an augmented generative system by explicitly fine-tuning a discriminator or prompting LLMs to improve contradiction discrimination.
3.2 Misinformation Reasoning
Erroneous instructions mislead model outputs more severely than inconsistent ones, as they lack obvious contradictions, requiring the model to comprehend, reason, compare input knowledge with its parameterized knowledge, and make objective judgments Cheang et al. (2023); Xu et al. (2024a). From the input perspective, this paper classifies the sources of erroneous information into two categories as follows: 1) Temporal Alignment Failure (Figure 4 B.1). arises when the knowledge provided by the user and the model is temporally misaligned due to updates occurring at different times, leading to inconsistent responses. Such discrepancies typically originate during the training process. 2) Information Contamination (Figure 4 B.2). refers to the degradation of model quality caused by the intentional distortion of input data.
To solve the above problems, existing methods mainly focus on improving model susceptibility in the face of internal and external knowledge conflicts through targeted fine-tuning and processing. CKL Jang et al. (2022) ensures that the model’s knowledge is updated in a timely manner through an online approach, although this approach is slightly weaker than re-training in terms of effectiveness Liska et al. (2022). BIPIA Yi et al. (2023) used adversarial training to combat the effects of information pollution and improve model robustness. CAR Weller et al. (2024) achieved nearly 20% improvement by discriminating external knowledge that may not be contaminated in the RAG system.
3.3 Fuzzy Language Interpretation
When user instructions contain fuzzy terms (e.g., polysemy or vagueness), LLMs may select an incorrect interpretation from multiple possibilities, potentially leading to misleading responses. This phenomenon has been observed across multiple information-seeking tasks Kim et al. (2024), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Self-defined problem (Figure 4 C.1). When the user inputs content with fuzzy sentences, LLMs may choose to generate content based on the preferences of its own training data. 2) Select data based on fuzzy input (Figure 4 C.2). In response to ambiguous user input, LLMs may select a default explanation without actively asking the user to clarify.
To solve the problem, researchers have started to experiment with cue engineering and relation alignment. Folkscope Yu et al. (2023) proposed the FolkScope framework, which uses a large language model to analyze and discriminate users’ fuzzy purchasing intention. Miko Lu et al. (2024) introduces a multimodal hierarchical intention generation framework that interprets users’ posting behaviors by analyzing the fuzzy information they share on social platforms. ATC Deng et al. (2023) utilizes the Active Thinking Chain cueing scheme, which enhances the proactivity of a biglanguage model by adding goal-planning capabilities to the descriptive reasoning chain.
3.4 Intention Clarification Failure
Unlike humans, who reason based on experience, LLMs lack real-world common sense and thus struggle to infer complex contexts beyond their input data. Moreover, they often fail to maintain a consistent reasoning trajectory across long texts, complex contexts, or multi-turn conversations. When handling intricate intentions or sentiment shifts, LLMs may struggle to retain prior context, leading to errors in inferring implicit user needs. We categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fails to detect sarcasm (Figure 4 D.1), when LLMs fails to understand the sarcastic intention of the user. 2) Ignores prior emotional context (Figure 4 D.2), when LLMs focus only on the second half of the sentence and ignore the emotions in the previous round of dialog.
To solve above problems, researchers have started to try to construct multi-domain datasets containing implicit intentions to strengthen the ability of the LLMs to reason about complex intentions and user emotions in multi-round interaction scenarios. DeepSeek-R1 Guo et al. (2025) enhances its understanding of human intention through a structured process with two RL stages for refining reasoning patterns and aligning with human preferences. S1 Muennighoff et al. (2025) uses budget forcing to control the number of thinking tokens. The upper limit is terminated early by a delimiter, while the lower limit prohibits delimiters and adds ”wait” to guide in-depth reasoning and optimize the quality of the answer. SoulChat Chen et al. (2023) fine-tuned LLMs by constructing a dataset containing more than 2 million samples of multi-round empathic conversations. MoChat Mo et al. (2024) constructs multi-round dialogues for spatial localization by using joint grouping spatio-temporal localization.
4 Realiable Generation
Despite strong performance, LLMs struggle with output reliability. Trained on large corpora using maximum likelihood estimation, they generate deterministic responses. While effective on familiar data, they often produce unstable or incomplete responses to unseen inputs, undermining reliability. We analyze LLM response stability and alignment in depth.
4.1 Response Stability
The knowledge acquired by the LLMs is generally determined in the pre-training stage and stored in a parameterized form. For data in a specific field, the current model is generally optimized by fine-tuning instructions so that it outputs what humans want Radford et al. (2019). If knowledge samples that the model has not seen are used in instruction-tuning, it will inevitably cause the model to give a definite response to unknown inputs, and there is a high probability that an answer will be fabricated. This is the over-confidence of the model that causes the model to output unreliable answers, and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fabricated incorrect information (Figure 5 A.1). When the model’s knowledge did not match the input question, it fabricated information that did not match the facts. 2) Incorrect context output (Figure 5 A.2). When the model’s knowledge did match the input question, it output incorrect context information. To address these issues, researchers have explored uncertainty, which quantifies the credibility and stability of model outputs as follows.
Figure 5: Common generation issues in LLMs include: (A) Unstable Content Generation (§4.1), where the model fabricates details when it lacks relevant knowledge (A.1) or produces incorrect contextual information despite having relevant knowledge (A.2). (B) Misalignment with Human Values (§4.2), where the model generates harmful or offensive content (B.1) or provides responses that conflict with moral and ethical standards (B.2).
Fine-tuning LLMs.
To make LLMs more accurate in estimating uncertainty, existing methods fine-tune models Sensoy et al. (2018); Amini et al. (2020). LUQ Zhang et al. (2024a) is a novel sampling-based uncertainty quantification method specifically designed for long texts. ConformalFactuality Mohri and Hashimoto (2024) defines the associated uncertainties for each possible output.
External Tools.
Fine-tuning LLMs typically demands substantial computing resources and slow training; therefore, reducing computational overhead is crucial for improving efficiency. Researcher has proposed methods to evaluate the uncertainty of model outputs through external tools Liu et al. (2024). CalibrateMath Lin et al. (2022) assesses uncertainty by requiring models to generate numerical answers with confidence levels, evaluating their reliability.
4.2 Alignment
Despite the impressive capabilities of large language models (LLMs), they have raised significant concerns regarding the unsafe or harmful content they may generate. LLMs are typically trained on vast datasets scraped from the internet, including inappropriate or harmful content Bender et al. (2021). This means that the models may inadvertently produce outputs misaligned with human values as follows: 1) Generation of Toxic Content (Figure 5 B.1). LLMs may generate toxic content, such as hate speech or offensive comments, when asked to respond to sensitive topics Luong et al. (2024); Dutta et al. (2024). 2) Conflicts with Moral/Ethical Standards (Figure 5 B.2). LLMs might produce outputs that conflict with moral or ethical standards, such as guiding illegal activities Ramezani and Xu (2023); Abdulhai et al. (2024). To tackle the previously mentioned concerns regarding unsafe or harmful content produced by LLMs, researchers have focused on various stages:
Pretraining Data Cleaning and Curation.
To minimize the risks associated with harmful or inappropriate content, LLM training datasets should undergo rigorous cleaning processes Bender et al. (2021), such as filtering out toxic language, hate speech, and harmful stereotypes. Tools like word embedding debiasing methods Rakshit et al. can help identify and remove toxic and biased content.
Table 1: A selection of widely used benchmark datasets for evaluating LLMs. The ’Type’ column categorizes the benchmarks based on task types, while ’Lang.’ indicates the language of the dataset, and ’Citation’ denotes the number of citations each benchmark has received.
Type Benchmark Year Lang. Description Citation
[…]
MT-BENCH-101 Bai et al. (2024) 2024 En Evaluation of multi-turn dialogue capability in LLMs. 39
LLM Reasoning BIPIA Yi et al. (2023) 2023 En Assesses LLMs vulnerability to hint injection attacks. 60
Miko Lu et al. (2024) 2024 En Evaluates multimodal LLMs in understanding social intention. 2
CONTRADOC Li et al. (2024a) 2023 En Detects self-contradictions in long documents. 7
CDCONV Zheng et al. (2022) 2022 Zh Detects contradictions in Chinese dialogues. 12
LLM Generation Open-LLM-Leaderboard Ye et al. (2024) 2024 En Assesses LLMs uncertainty in generated content. 19
ETHICS Dataset Hendrycks et al. (2021) 2020 En Evaluates LLMs on moral reasoning and ethics. 468
FACTOR Muhlgay et al. (2024) 2023 En Measures factual accuracy in LLM-generated text. 71
Reinforcement Learning-based Alignment.
To further align LLMs with human values and societal norms, reinforcement learning methods, such as RLHF and its advanced variants, such as PPO Ouyang et al. (2022), DPO Zeng et al. (2024), and GRPO Shao et al. (2024), are very essential. As an extension of RLHF, RLAIF Lee et al. (2024) leverages AI systems to assist in the feedback process, making evaluation and fine-tuning more scalable.
In-context Alignment.
In-context alignment leverages the ability of LLMs to adapt their responses based on a few examples provided in the prompt. Lin et al. (2024) demonstrates that effective alignment can be achieved purely through ICL with just a few stylistic examples and a system prompt. Huang et al. (2024a) explored the effectiveness of different components of In-context alignment, and found that examples within the context are crucial for enhancing alignment capabilities.
5 Benchmark
This section covers benchmarks for LLMs in instruction understanding, reasoning, and generation (Table 1).
5.1 Benchmarking Instruction Understanding
Instruction understanding in LLMs involves extracting key information, maintaining coherence, and adapting to dynamic conversation changes, especially in long or multi-round dialogues. Zhang et al. Zhang et al. (2024c) provides a benchmark consisting of 12 tasks, with an average data length exceeding 100K tokens, designed to evaluate LLMs’ ability to handle long contexts. The results show that the performance of LLMs significantly declines when dealing with long contexts, suggesting further improvements. MT-BENCH-101 Bai et al. (2024) constructs a three-layer hierarchical capability classification and 1,388 dialogue pairs across 13 tasks, providing a comprehensive evaluation of LLMs’ fine-grained capabilities in multi-turn dialogue. Experimental results show that commonly used techniques do not significantly enhance multi-turn capabilities.
5.2 Benchmarking LLM Reasoning
LLM Reasoning involves inferring user intentions by interpreting both explicit and implicit language cues. BIPIA Yi et al. (2023) evaluates LLMs under indirect hint injection attacks across five scenarios and 250 targets, revealing vulnerabilities in all models, with GPT-3.5-turbo and GPT-4 exhibiting notably higher susceptibilities. Miko Lu et al. (2024) assesses multimodal models in understanding social media user intentions. The benchmark, which includes 979 social media entries, shows that multimodal LLMs outperform text-only models like LLama2-7B and GLM4. Incorporating image data enhances the model’s ability to interpret user intentions, improving accuracy. CONTRADOC Li et al. (2024a) is the first dataset for analyzing self-contradictions in long documents. Evaluation of GPT-4, PaLM2, and other LLMs on this dataset reveals that while GPT-4 outperforms others and even surpasses human performance, it still struggles with complex contradictions requiring nuanced reasoning. CDCONV Zheng et al. (2022) focuses on contradiction detection in Chinese dialogues, containing over 12,000 dialogue rounds. The study shows that the Hierarchical method consistently outperforms others in detecting contradictions, highlighting the importance of accurate contextual modeling in dialogue understanding.
5.3 Benchmarking LLM Generation
LLM Generation assesses a model’s ability to understand user instructions, avoid fabricating false information, and generate accurate, contextually appropriate responses. Open-LLM-Leaderboard Ye et al. (2024) introduces a novel benchmark that integrates uncertainty quantification to evaluate the reliability of content generation across tasks like QA, comprehension, and dialogue. Results show that larger models often exhibit greater uncertainty, and fine-tuned models tend to have higher uncertainty despite higher accuracy. ETHICS Hendrycks et al. (2021) evaluates whether generated content aligns with human ethical values, such as justice and well-being. The study finds that while models like GPT-3 show promise in predicting human moral judgments, they still need improvement in this domain. FACTOR Muhlgay et al. (2024) addresses the evaluation of factuality in LLMs by providing a scalable method that ensures diverse and rare facts are considered. Testing models such as GPT-2 and GPT-Neo show that, while benchmark scores correlate with perplexity, they better reflect factuality in open-ended generation, especially when retrieval augmentation is applied.
6 Future Directions
This section summarizes ongoing challenges in instruction understanding, reasoning, and reliable generation with LLMs and outlines potential future research directions.
Automated Annotation Framework.
Although LLMs excel in general-domain tasks, they often produce hallucinated or incomplete content in specialized fields due to limited domain-specific training data. While contextual learning and instruction fine-tuning methods have been explored to address this issue, manual data annotation remains labor-intensive and prone to quality inconsistencies. An automated annotation framework could streamline data labeling, enhancing model performance in specialized fields by ensuring higher quality and scalability of domain-specific training datasets.
GraphRAG.
LLMs have shown impressive language generation capabilities through pre-training on large datasets, but their reliance on static data often results in inaccurate or fictional content, particularly in domain-specific tasks. The Graph-enhanced generation approach aims to tackle this by leveraging KGs and GNNs for precise knowledge retrieval. Despite its advantages, GraphRAG faces challenges in capturing structural information during graph reasoning tasks and struggles with multi-hop retrieval accuracy and conflict resolution between external and internal knowledge. Future work should focus on refining retrieval strategies and improving the stability and accuracy of GraphRAG in complex tasks.
Quantifying Uncertainty in LLMs.
As LLMs like LLama and ChatGPT have revolutionized content generation, ensuring the reliability of their outputs remains a critical challenge. Uncertainty quantification is a promising approach to address this, enabling models to provide a confidence assessment alongside their responses. Evidence learning Sensoy et al. (2018); Amini et al. (2020), an emerging method in uncertainty representation, offers a reliable approach for quantifying uncertainty directly from data. However, its application to LLMs is computationally intensive, and most existing work focuses on small-scale models. Future research should aim to optimize uncertainty quantification for large-scale models efficiently.
Balancing Safety and Performance.
Although advancements in alignment techniques have improved factual accuracy and safety, they often come at the cost of the model’s creativity and fluency. Striking a balance between safety and performance is crucial. Future research should explore new alignment methods that ensure both the safety and usability of LLMs, optimizing the trade-off between generating reliable, safe content and maintaining the model’s creative and contextual capabilities.
7 Conclusion
This paper analyzes LLMs’ performance in processing user instructions. Despite progress in natural language understanding, LLMs struggle with complex, inconsistent instructions, often resulting in biases, errors, and hallucinations. Improvements through prompt engineering, knowledge retrieval, model expansion, fine-tuning, and RLHF have not fully addressed LLMs’ limitations in reasoning and comprehension, limiting their real-world applicability. We identify three key bottlenecks: instruction comprehension, reasoning ability, and reliable content generation. Future research should focus on enhancing reasoning for complex instructions and aligning outputs with user intent, to improve LLMs’ adaptability and reliability in real-world tasks.